# On the PM2.5 – Mortality Relationship: A Bayesian Dynamic Model for Spatio-Temporal Confounding

### v1

@@ E.2 Performance Assessment
To complete the results of Simulation Study 1, Table E.4 reports, for each scenario and model, the value for the three indices, namely:
@@
  • average mean squared error, $MSE_{avg}=\frac{1}{{T}}\sum_{t=1}^{{T}}\left[\frac{1}{100}\sum_{i=1}^{100}\left({\rm\mathbb{E}}[P_{t}|\bullet]-P_{t}\right)^{2}\right]$
  • maximum absolute bias, $MAB=\max_{t}|{\rm\mathbb{E}}[P_{t}|\bullet]-P_{t}|$
@@
  • average mean absolute error, $MAE_{avg}=\frac{1}{{T}}\sum_{t=1}^{{T}}\left[\frac{1}{100}\sum_{i=1}^{100}|{\rm\mathbb{E}}[P_{t}|\bullet]-P_{t}|\right]$
@@
Here, the differences are computed with respect to $P_{t}$, i.e. the true simulated percent change of risk.
Table E.4 indicates that, if the exposure effect is constant, the GLMadj model should be preferred, otherwise the DGLM is the best choice. Note, however, that the shape of the exposure effect is not known in practice, so the DGLM is the most robust choice across all scenarios.
@@
Table E.4: Simulation Study 1: $MSE_{avg}$, $MAB$, and $MAE_{avg}$ calculated for each model considered under each Scenario. The lowest values are indicated in italic.
@@
Scenario Exposure Effect DGLM Null GLMadj DGLM Null GLMadj DGLM Null GLMadj
S1 Constant 0.642 63.158 0.368 0.728 7.929 0.568 0.686 7.929 0.568
Periodic 0.697 115.672 3.276 0.833 12.678 3.109 0.717 10.643 1.490
Complex 0.672 66.741 1.283 1.217 9.912 2.423 0.691 8.099 0.989
S2 Constant 6.886 19.741 6.622 3.660 4.435 2.547 2.571 4.435 2.547
Periodic 7.177 32.742 10.620 3.337 7.566 4.935 2.625 5.531 2.900
Complex 6.782 20.939 8.094 3.691 6.285 4.476 2.547 4.472 2.664
S3 Constant 0.251 45.706 0.085 0.406 6.760 0.260 0.331 6.760 0.260
Periodic 0.276 86.516 2.706 0.480 11.224 2.823 0.360 9.189 1.396
Complex 0.271 48.538 0.981 0.911 8.717 2.126 0.325 6.904 0.852
E.3 On The Removal of Large-Scale Temporal Trend
@@
We argue that the term $X_{it}-\widehat{X_{t}}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$, should be used in place of simply $X_{it}$ in the DGLM. Removing the overall temporal trend from the exposure not only follows from theoretical arguments (see Equation (3) in the main paper), but also it is crucial to control for confounding bias. The local variations in the exposure are in fact less likely to be confounded by the unmeasured confounder, which is assumed to be spatially and temporally smooth (Janes et al., 2007, Greven et al., 2011) – see also Section 3 of the manuscript. To illustrate this with simulated data, we compare the proposed DGLM with a reparameterized model that does not remove the overall trend from the exposure. Essentially, Equation (6) in the paper can be rewritten as:
$\vartheta_{it}=o_{it}+\beta_{0it}^{*}+\beta_{1t}X_{it}+{{}^{m}\mathbf{Z}}_{it}^{\prime}\bm{\alpha}_{m}+\tilde{u}_{it}\,,$
where
$\beta_{0it}^{*}=\beta_{0it}-\beta_{1t}\widehat{X_{t}}\,.$
@@
We refer to this model as DGLM_single, since it is a single-step approach that does not require to estimate the overall temporal trend.
The results, in terms of percent change of risk associated with a 10-unit increase in the exposure, $P_{t}$, for each scenario considered in the main paper are depicted in Figure E.1. The estimates tend to differ at the beginning and at the end of the time window, where the DGLM_single model is more biased. In addition, note that the DGLM_single model is not able to recover the constant effect in the top-left subplot. For completeness, Table E.5 reports the $MSE_{avg}$, $MAB$, and $MAE_{avg}$ for both model considered under each scenario, where it is confirmed that the proposed DGLM outperforms the DGLM_single model in all cases.
Figure E.1: Simulation Study 1: results are expressed in terms of percent change of risk associated with a 10-unit increase in the exposure, $P_{t}$. The black line represents the true exposure effect (whose shape differs in each row). The solid red and dashed brown lines represent the average of the posterior means of $P_{t}$ obtained from DGLM and DGLM_no over 100 replicates, respectively. The shaded areas represent the percentiles $2.5$ and $97.5$ of the posterior means of $P_{t}$ under each approach.
Table E.5: Simulation Study 1: $MSE_{avg}$, $MAB$, and $MAE_{avg}$ calculated for the DGLM and DGLM_single models under each Scenario. The lowest values for each index are indicated in italic.
Exposure $MSE_{avg}$ $MAB$ $MAE_{avg}$
Scenario Effect DGLM DGLM_single DGLM DGLM_single DGLM DGLM_single
S1 Constant 0.642 0.873 0.728 2.002 0.686 0.773
Periodic 0.697 0.816 0.833 1.762 0.717 0.763
Complex 0.672 0.890 1.217 1.955 0.691 0.799
S2 Constant 6.886 7.711 3.660 4.351 2.571 2.707
Periodic 7.177 7.695 3.337 4.178 2.625 2.706
Complex 6.782 7.454 3.691 4.333 2.547 2.660
S3 Constant 0.251 0.254 0.406 1.033 0.406 0.418
Periodic 0.276 0.250 0.480 0.957 0.428 0.417
Complex 0.271 0.282 0.911 1.183 0.418 0.436
Appendix F Simulation Study 2
We now illustrate Simulation Study 2, where we want to assess whether the proposed DGLM is able to recover the first derivative of the unknown functional form, which represents the marginal effect of the exposure on the outcome. To this end, we consider Scenarios S1 and S2 again, but we assume that the linear predictor is as follows:
$\vartheta_{it}=o_{it}+\beta_{0i}+f(X_{it};\bm{\gamma}_{f})+Z_{it}+u_{it}\,,$
where $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}+\gamma_{f2}X_{it}^{2}$, $\gamma_{f0}=\gamma_{f1}=0$, $\gamma_{f2}=2.5\times 10^{-5}$. Henceforth, the true values of the time-varying coefficients are $\beta_{1t}=5\times 10^{-5}\widehat{X_{t}}$, and $\beta_{0it}=\beta_{0i}+2.5\times 10^{-5}\widehat{X_{t}}^{2}+Z_{it}$. All the other parameters are the same as in the previous Section.
Scenario S1.
Figure F.2 refers to Scenario S1 and consists of two subplots. The top subplot compares spatial averages of the time-varying intercepts. That is,
$\overline{\beta_{0t}}=\frac{1}{{N}}\sum_{i=1}^{{N}}\beta_{0it}\quad\mbox{for}\quad t=1,\dots,{T}$
and
$\overline{{\rm\mathbb{E}}[\beta_{0t}|\bullet]}=\frac{1}{{N}}\sum_{i=1}^{{N}}{\rm\mathbb{E}}[\beta_{0it}|\bullet]\quad\mbox{for}\quad t=1,\dots,{T}$
denote the average true intercept and average posterior mean, respectively. These quantities are different across the 100 replicates, so the blue (green) line corresponds to the average $\overline{\beta_{0t}}$ ($\overline{{\rm\mathbb{E}}[\beta_{0t}|\bullet]}$), while the shaded area of the same color defines the percentiles $2.5$ and $97.5$. The blue and green shaded areas do not completely overlap, hence the unmeasured confounder can be captured only in part. The bottom subplot shows that the true value of $\beta_{1t}$ (black line) is almost always within the red shaded area, indicating that confounding bias is alleviated but not completely ruled out.
Figure F.2: Simulation Study 2: results from Scenario S1 with $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}+\gamma_{f2}X_{it}^{2}$. Top: the blue solid line and the blue shaded area are the average and the percentiles $2.5$ and $97.5$ of the spatial mean of the simulated intercepts over 100 replicated datasets; the green solid line and green shaded area depict the same quantities but for the corresponding spatially-averaged posterior mean obtained from the proposed DGLM. Bottom: the black solid line represents the true exposure effect, while the red solid line and the shaded area represent the average and the percentiles $2.5$ and $97.5$ of the posterior means of $\beta_{1t}$.
Scenario S2.
As for Scenario S2, Figure F.3 indicates that it is not possible to completely recover the true function, $f(\cdot;\cdot)$. The unmeasured confounder has low temporal autocorrelation, and the posterior means of the intercepts tend to have a rougher behavior. However, some variation in the unmeasured confounder is “absorbed” by the local exposure effect, which remains biased.
Figure F.3: Simulation Study 2: results from Scenario S2 with $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}+\gamma_{f2}X_{it}^{2}$. Top: the blue solid line and the blue shaded area are the average and the percentiles $2.5$ and $97.5$ of the spatial mean of the simulated intercepts over 100 replicated datasets; the green solid line and green shaded area depict the same quantities but for the corresponding spatially-averaged posterior mean obtained from the proposed DGLM. Bottom: the black solid line represents the true exposure effect, while the red solid line and the shaded area represent the average and the percentiles $2.5$ and $97.5$ of the posterior means of $\beta_{1t}$.
Scenario without confounding.
To understand whether the residual bias in Figures F.2 and F.3 is due to the Taylor’s approximation, we now assume that the outcome is not affected by any confounder, i.e. ${\bf Z}={\bf 0}$. We only report the results for the Scenario S1 where $\phi_{x}^{S}=\phi_{x}^{T}=0.2$. The estimated exposure effect is depicted in Figure F.4, which indicates that the proposed DGLM is able to recover the true effect, with no visible difference between the black and red lines for almost all the temporal domain.
Figure F.4: Simulation Study 2: results from Scenario S1 without confounding. The black solid line represents the true exposure effect. The red solid line and the shaded area represent the average and the percentiles $2.5$ and $97.5$ of the posterior means of $\beta_{1t}$ obtained from the proposed DGLM over 100 replicated datasets.
Appendix G Application: Comparison with Restricted Model
@@
We assess the impact that imposing some constraints on the proposed DGLM has on the estimated exposure effect. Specifically, we remove the measured confounders from the model specification. This is equivalent to imposing that $\bm{\alpha}_{m}={\bf 0}$ in Equation (6) of the main text. Hence, the linear predictor and the temporal dynamics of the coefficients are specified as follows:
$\bm{\vartheta}_{t}$ $={\bf o}_{t}+\mathbf{\tilde{X}}_{t}\tilde{\bm{\beta}}_{t}+\tilde{{\bf u}}_{t}\,,$
$\tilde{\bm{\beta}}_{t}$ $=\tilde{\bm{\beta}}_{t-1}+\mathbf{w}_{t}\,.$
@@
We refer to the implied model as DGLM-X to emphasize that the exposure is the only covariate. Figure G.5 shows the posterior mean of the percent change of risk, as defined above, under the DGLM-X (in green) and the proposed DGLM (in red). It suggests that, in general, the temporal dynamic of the exposure effect would not be affected much if the measured confounders were not included in the model. Henceforth, the proposed DGLM can provide accurate estimates even in complex scenarios.
@@
However, it is worth noting that the DGLM-X provides greater estimates than the proposed estimates during the summer, while during the winter the estimates are quite similar. This suggests that the confounding effect may vary across time. We hypothesize that this is due to air temperature that exceeds a certain threshold; in other words, confounding may occur due to heat waves. In fact, heat waves are demonstrated to have an increased effect on mortality during summer days with relatively high levels of PM2.5 concentrations, and this positive correlation may be the source of the overestimation (Analitis et al., 2008, 2014). Another study observed a similar result for the population aged 75 years or more (Analitis et al., 2018). Another possible explanation for the overestimation could be the modification of the PM2.5 chemical composition, with high levels of the most toxic compounds developing in summer (Vinson et al., 2020).
@@
Figure G.5: Estimated time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The red line represents the posterior mean of the proposed DGLM, while the green line the posterior mean of the DGLM-X.

### v2

@@ E.2 Performance Assessment
To complete the results of Simulation Study 1, Tables E.4E.6 report, for each scenario and model, the value for the three indices, namely:
@@
  • average mean squared error, $MSE_{avg}=\frac{1}{{N}{T}}\sum_{i=1}^{{N}}\sum_{t=1}^{{T}}\left[\frac{1}{100}\sum_{j=1}^{100}\left({\rm\mathbb{E}}^{(j)}[P_{it}|\bullet]-P_{it}\right)^{2}\right]$, where ${\rm\mathbb{E}}^{(j)}[P_{it}|\bullet]$ is the posterior mean for the $j$-th simulated dataset;
  • maximum absolute bias, $MAB=\max_{i,t}\frac{1}{100}\sum_{j=1}^{100}|{\rm\mathbb{E}}^{(j)}[P_{it}|\bullet]-P_{it}|$
@@
  • average mean absolute error, $MAE_{avg}=\frac{1}{{N}{T}}\sum_{i=1}^{{N}}\sum_{t=1}^{{T}}\left[\frac{1}{100}\sum_{i=1}^{100}|{\rm\mathbb{E}}^{(j)}[P_{it}|\bullet]-P_{it}|\right]$
@@
Here, the differences are computed with respect to $P_{it}=100(\exp\{10(\beta_{1it})\}-1)$, i.e. the true simulated percent change of risk, which varies spatio-temporally.
Tables E.4E.6 indicate that, SDGLMC is the best model among those compared in terms of $MSE_{avg}$, $MAB$, and $MAE_{avg}$ for all scenarios. Recalling the decomposition of $\beta_{1it}$ in Section 3.1 of the main manuscript, we have also computed the average mean squared error for each of the components of the exposure effect – global, temporal and spatial. SDGLMC provides the lowest values in most of the cases, and the improvements over the other approaches are more pronounced for the temporal component.
Since we want the overall error to be as low as possible, and since the shape of the exposure effect is not known in practice, simulation results suggest that the proposed model is the most robust choice across all scenarios.
@@
Table E.4: Simulation Study 1: $MSE_{avg}$ calculated for each model considered under each Scenario. The lowest values are indicated in italic.
Scenario Component Formula SDGLMC Null GLMadj Dummy Periodic JDZ
S1 Overall $100(\exp\{10\beta_{1it}\}-1)$ 0.669 36.214 2.417 1.888 2.780 2.522
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.115 34.568 0.193 0.142 0.225 0.275
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.200 1.418 1.418 0.889 1.705 1.418
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.354 0.228 0.806 0.855 0.844 0.829
S2 Overall $100(\exp\{10\beta_{1it}\}-1)$ 5.617 20.440 6.596 6.022 8.118 6.681
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 5.246 13.331 4.807 4.897 6.568 4.893
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.103 1.418 1.418 0.837 1.204 1.418
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.268 5.691 0.371 0.277 0.288 0.369
S3 Overall $100(\exp\{10\beta_{1it}\}-1)$ 0.857 31.247 2.286 1.699 2.222 2.337
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.509 29.444 0.370 0.310 0.423 0.417
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.133 1.418 1.418 0.886 1.281 1.418
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.216 0.385 0.498 0.500 0.506 0.502
Table E.5: Simulation Study 1: $MAB$ calculated for each model considered under each Scenario. The lowest values are indicated in italic.
Scenario Component Formula SDGLMC Null GLMadj Dummy Periodic JDZ
S1 Overall $100(\exp\{10\beta_{1it}\}-1)$ 1.212 7.770 3.816 3.796 4.010 3.618
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.295 5.856 0.189 0.022 0.340 0.316
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.276 2.100 2.100 1.913 2.497 2.100
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.934 1.373 1.859 1.873 1.787 1.788
S2 Overall $100(\exp\{10\beta_{1it}\}-1)$ 2.898 5.896 3.958 4.048 5.117 3.904
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 2.280 3.647 2.183 2.204 2.555 2.203
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.573 2.100 2.100 2.003 2.190 2.100
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 1.198 1.095 0.387 0.356 0.372 0.313
S3 Overall $100(\exp\{10\beta_{1it}\}-1)$ 1.506 7.683 2.767 2.769 3.607 2.791
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.599 5.413 0.479 0.403 0.529 0.523
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.211 2.100 2.100 1.968 2.250 2.100
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.740 1.691 1.100 1.024 1.114 1.168
Table E.6: Simulation Study 1: $MAE_{avg}$ calculated for each model considered under each Scenario. The lowest values are indicated in italic.
Scenario Component Formula SDGLMC Null GLMadj Dummy Periodic JDZ
S1 Overall $100(\exp\{10\beta_{1it}\}-1)$ 0.654 5.856 1.287 1.111 1.330 1.315
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.297 5.856 0.340 0.286 0.400 0.422
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.355 1.065 1.065 0.801 1.046 1.065
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.468 0.374 0.728 0.751 0.744 0.739
S2 Overall $100(\exp\{10\beta_{1it}\}-1)$ 2.281 3.868 2.226 2.213 2.564 2.242
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 2.280 3.647 2.183 2.204 2.555 2.203
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.253 1.065 1.065 0.794 0.823 1.065
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.402 1.908 0.485 0.415 0.425 0.485
S3 Overall $100(\exp\{10\beta_{1it}\}-1)$ 0.735 5.413 1.272 1.060 1.176 1.285
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.599 5.413 0.479 0.404 0.529 0.523
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.287 1.065 1.065 0.806 0.907 1.065
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.365 0.495 0.557 0.556 0.561 0.558
Appendix F Simulation Study 2
We now illustrate Simulation Study 2, where we want to assess whether SDGLMC is able to recover the first derivative of the unknown functional form, which represents the marginal effect of the exposure on the outcome. To this end, we consider the Scenario S1 again, but we assume that the linear predictor is as follows:
$\vartheta_{it}=o_{it}+\beta_{0}+f(X_{it};\bm{\gamma}_{f})+Z_{it}+u_{it}\,,$
where $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}+\gamma_{f2}X_{it}^{2}+\gamma_{f3}X_{it}^{3}$, $\gamma_{f0}=-0.08$, $\gamma_{f1}=0.02$, $\gamma_{f2}=-8\times 10^{-5}$, $\gamma_{f3}=5\times 10^{-8}$. Henceforth, the true values of the time-varying coefficients are $\beta_{1t}=0.02-1.6\times 10^{-4}\widehat{X}_{it}+1.5\times 10^{-7}\widehat{X}_{it}^{2}$, and $\beta_{0it}=\beta_{0}-0.08+0.02\widehat{X}_{it}-8\times 10^{-5}\widehat{X}_{it}^{2}+5\times 10^{-8}\widehat{X}_{it}^{3}+Z_{it}$. All the other parameters are the same as in the previous Section.
Table F.7 reports the $MSE_{avg}$, $MAB$, and $MAE_{avg}$ that were obtained. The small values of the indices indicate that the proposed model is able to recover the first derivative of the unknown functional form. Some confounding bias is still present, but it is not attributable to the Taylor’s series approximation, since it is not present when the data are generated without unmeasured confounder (i.e., $Z_{it}=0$).
Table F.7: Simulation Study 2: $MSE_{avg}$, $MAB$, and $MAE_{avg}$ calculated for SDGLMC and SDGLMC_single under Scenario S1 with $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}+\gamma_{f2}X_{it}^{2}+\gamma_{f3}X_{it}^{3}$. The lowest values for each index are indicated in italic.
@@ Appendix F Simulation Study 2
Component SDGLMC SDGLMC_single SDGLMC SDGLMC_single SDGLMC SDGLMC_single
Overall 1.381 1.897 2.288 3.330 0.929 1.094
Global 0.050 0.398 0.170 0.614 0.185 0.614
Temporal 0.376 0.503 2.000 2.579 0.481 0.544
Spatial 0.945 0.984 0.452 0.860 0.762 0.741
F.1 On The Removal of Large-Scale Temporal Trend.
@@
We argue that the term $X_{it}-\widehat{X}_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$, should be used in place of simply $X_{it}$ in the DGLM. Removing the overall temporal trend from the exposure not only follows from theoretical arguments (see Equation (3) in the main paper), but also it is crucial to control for confounding bias. The local variations in the exposure are in fact less likely to be confounded by the unmeasured confounder, which is assumed to be spatially and temporally smooth (Janes et al., 2007, Greven et al., 2011) – see also Section 3 of the manuscript. To illustrate this with simulated data, we compare the proposed SDGLMC with a reparameterized model that does not remove the overall trend from the exposure. Essentially, if the first-order Taylor’s series is evaluated at 00, Equation (6) in the paper can be rewritten as:
${\color[rgb]{0,0,0}\vartheta_{it}=o_{it}+\beta_{0it}^{*}+\beta_{1t}^{*}X_{it}+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+\tilde{u}_{it}\,,}$
where $\beta_{1it}^{*}=f^{\prime}(0;\bm{\gamma}_{f})$ and $\beta_{0it}^{*}=\beta_{0}+f(0;\bm{\gamma}_{f})+Z_{it}$.
@@
We refer to this model as SDGLMC_single, since it is a single-step approach that does not require to estimate the overall spatio-temporal trend.
The results for the reparameterized approach are reported in Table F.7, where it is confirmed that SDGLMC outperforms SDGLMC_single according to all indices.
Appendix G Application: Further Analyses
G.1 Comparison with Restricted Model
@@
We assess the impact that imposing some constraints on the proposed approach has on the estimated exposure effect. Specifically, we remove the measured confounders from the model specification. This is equivalent to imposing that $\bm{\alpha}={\bf 0}$ in Equation (7) of the main text. Hence, the linear predictor is specified as follows:
${\color[rgb]{0,0,0}\bm{\vartheta}_{t}=\mathbf{o}_{t}+{\bf 1}_{N}(\overline{\delta}_{0}+\delta_{0t})+\tilde{\bm{\delta}}_{0}+\bm{\delta}_{0t}^{*}+\mathbf{\tilde{X}}_{t}(\overline{\delta}_{1}+\delta_{1t})+\mathbf{\tilde{X}}_{t}\odot\tilde{\bm{\delta}}_{1}+\tilde{{\bf u}}_{t}\,.}$
@@
We refer to the implied model as SDGLMCX to emphasize that the exposure is the only covariate. Figure G.1 shows the posterior mean of the percent change of risk, as defined above, under the SDGLMCX (in green) and the proposed DGLM (in red). It suggests that, in general, the temporal dynamic of the exposure effect would not be affected much if the measured confounders were not included in the model. Henceforth, the proposed approach can provide accurate estimates even in complex scenarios.
@@
However, it is worth noting that SDGLMCX provides greater estimates than SDGLMC during the summer, while during the winter the estimates are quite similar. This suggests that the confounding effect may vary across time. We hypothesize that this is due to air temperature exceeding a certain threshold; in other words, confounding may occur due to heat waves. In fact, heat waves are demonstrated to have an increased effect on mortality during summer days with relatively high levels of PM2.5 concentrations, and this positive correlation may be the source of the overestimation (Analitis et al., 2008, 2014). Another study observed a similar result for the population aged 75 years or more (Analitis et al., 2018). Another possible explanation for the overestimation could be the modification of the PM2.5 chemical composition, with high levels of the most toxic compounds developing in summer (Vinson et al., 2020).
@@
Figure G.1: Estimated time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The red line represents the posterior mean of SDGLMC, while the green line the posterior mean of SDGLMCX.
G.2 Comparison with a GLM with interactions
We argue that the summer peaks of the exposure effect are likely due to potential interactions between PM2.5 and temperature or other confounders. To provide further evidence, we compare the proposed approach with a model that includes an interaction term between PM2.5and each measured confounder. Therefore, we consider a hierarchical model, referred to as GLMint, that follows the same structure as defined in Section D, but with $\mathbf{A}_{it}=(X_{it},X_{it}{{\bf M}}_{it}^{\prime})^{\prime}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$.
Given the hierarchical structure of GLMint, the exposure effect can be estimated for each HD by summing the PM2.5 coefficient and the interaction coefficients (multiplied by the confounder values measured in the respective HD at each point in time). The estimates are then averaged across the HDs to obtain a single estimate for each time point.
Figure G.2 shows the posterior mean of the percent change of risk (as defined in previous sections) under SDGLMC (in red) and GLMint (in dark blue). The estimates from both models appear similar during winter. However, GLMint may overestimate the exposure effect during summer, with peaks approximately equal to $16\%$, which is unexpected for short-term associations. This suggests that GLMint is not able to capture the temporal dynamics of the exposure effect as well as SDGLMC.
Given that the effect estimated by GLMint is not smooth in time, another hierarchical model, GLMint_smooth, is added to the comparison. It includes interaction terms between PM2.5 and smoothed versions of temperature and relative humidity, and is depicted in light blue in Figure G.2. The smoothed confounders are obtained by using a principal spline with $15$ spatial bases and $15$ temporal bases. The results show that the temporal dynamics under GLMint_smooth and SDGLMC are similar in 2018 and 2022. However, it seems that GLMint_smooth could underestimate the effect in 2019, and could overestimate it during the pandemic. Furthemore, GLMint_smooth and the Periodic model seem to provide similar estimates (see Figure 3 in the main manuscript).
Finally, this analysis corroborates the hypothesis of effect modification by unmeasured spatio-temporal confounders, which is not well captured by GLMint or GLMint_smooth. The results indicate that the possibility of a more complex interaction between PM2.5 and the confounders cannot be ruled out.
Figure G.2: Estimated time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations.